sir pls solve ths qutnthe circle x2+y2-6x-10y+c=0 does not touch or intersect the co-ordinate axes and the point(1,4) is inside the circle.find the set of values of c.

Grade:11

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear rohan

x2+y2-6x-10y+c=0

(x-3)2 +(y-5)2 = 34-c

circle does not intersect the co ordinate axis so √(34-c) <3

  or 34-c<9

 or  c >25

 

and 1,4 lie inside the circle so redius of circle must be greater then distance between the point (1,4) and (3,5)

√(34-c)> √(3-1)2 +(5-4)2

34-c  >√5

c < 34- √5

so   25 < c  < 34- √5


 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
We are all IITians and here to help you in your IIT JEE & AIEEE preparation.

All the best.

Regards,
Askiitians Experts
Badiuddin

 

sid
7 years ago
But sir ,answer that is mentioned here in book is (25,29).….......…....................///.........................